OpenStudy (anonymous):

A square pyramid has a base area of 25 ft2. The pyramid’s total surface area is 105 ft2. What is the area of each triangular face? A. 20 ft2 B. 26 ft2 C. 40 ft2 D. 80 ft2

OpenStudy (anonymous):

u find surface area by adding all the sides up

OpenStudy (cwrw238):

and each triangular face will have the same area

OpenStudy (anonymous):

 so in this case u would divide 105 by 4

OpenStudy (anonymous):

which equals 26.25, which means it is B?

OpenStudy (cwrw238):

no thats not correct

OpenStudy (cwrw238):

the total surface area includes the area of the base

OpenStudy (anonymous):

ok sorry i forgot that

OpenStudy (cwrw238):

so the area of the 4 triangles = 105-25 = 80 so now its easy to find the area of 1 triangle

OpenStudy (cwrw238):

* each triangle

OpenStudy (anonymous):

so it is A. 20?

OpenStudy (cwrw238):

80 / 4 = 20 yes

OpenStudy (anonymous):

Okay, thanks :)